# From Static to Adaptive Defense: Federated Multi-Agent Deep Reinforcement Learning-Driven Moving Target Defense Against DoS Attacks in UAV Swarm Networks

### v2

@@ V Performance Evaluation
Federated aggregation interval $K$ 20
@@
V-A Simulation Setup
The key hyperparameters, including the reward weighting coefficients $(\alpha,\beta,\zeta,\eta,\xi)$ and the entropy coefficient $\mu$, were determined through an iterative, empirical tuning process. This process aimed to find a balance between competing objectives, such as prioritizing attack mitigation over defense cost and managing the exploration-exploitation trade-off, to ensure the policy converges to a stable and effective defense strategy. The final values for all simulation parameters are summarized in Table I with reference to [38, 39, 40].
@@ V-C Effectiveness and Efficiency Analysis
To further analyze the defense effectiveness in this work, we measured the attack mitigation rate of proposed method PG-FMADRL under the previously defined attack scenarios and compare it with four representative MTD-based approaches including WF-MTD, RE-MTD, ID-HAM, DESOLATER.
It can be seen that PG-FMADRL consistently achieves the highest or near-highest attack mitigation rates across all attack scenarios. For node attack scenario, PG-FMADRL achieves highest attack mitigation rates of 0.9975 and 0.9999 against fixed and random attackers, respectively, while also maintaining the lowest standard deviation among all compared methods. This indicates superior stability and robustness. In the link attack scenario, although the attack mitigation rates show a slight decline, the proposed method still achieves the highest values among all methods, reaching 0.9969 for fixed and 0.9782 for random attackers. These results demonstrate the method’s strong and consistent defense effectiveness across different scenarios.
Notably, PG-FMADRL’s advantage becomes most apparent against sophisticated greedy attackers. While ID-HAM and DESOLATER rely on independent learning where each agent must discover effective strategies in isolation, our federated architecture acts as a knowledge accelerator. For instance, in the challenging greedy link attack scenario, PG-FMADRL reaches an attack mitigation rate of 0.9367, significantly outperforming the other methods, which achieve only 0.7910, 0.7625, 0.7360, and 0.7721, respectively. This is because our reward-weighted aggregation mechanism allows the entire swarm to rapidly learn from the experiences of the most successful agents those achieving high rewards, by effectively countering the attack. Our collaborative learning ensures that effective defense policies are quickly propagated and adopted across the swarm, leading to a more robust and coordinated response.
Beyond defense effectiveness, we also assess the efficiency of each method using recovery time as an indicator, which is defined as the duration required for the system to restore normal connectivity following an attack-induced disruption. Fig. 4 presents the recovery times of different defense methods.
As we can see, in the node attack scenario, WF-MTD shows the longest recovery times under fixed and random attackers, reaching 6.2 and 7.6 seconds, respectively. Under greedy attacks, RE-MTD performs the worst, with a recovery time of 18.7 seconds. In contrast, our PG-FMADRL method dramatically reduces the recovery time by approximately 94.6%, requiring only about 1 second to restore a secure state. In the link attack scenario, the worst recovery times are observed for different methods depending on the attacker type: ID-HAM for fixed attackers with 4.1 seconds, WF-MTD for random attackers with 3.7 seconds, and RE-MTD for greedy attackers with 14.7 seconds. Such prolonged recovery durations leave the UAV swarm vulnerable for extended periods, increasing both operational risk and energy consumption. Although DESOLATER performs relatively well under greedy attacks in the node scenario, its recovery time increases significantly to 11.8 seconds in the link scenario. On the other hand, PG-FMADRL achieves the shortest recovery times across all attack scenarios and attacker types, enabling rapid mitigation and minimizing disruption.
In summary, the superior attack mitigation rates and reduced recovery times of PG-FMADRL stem from its reward-weighted aggregation and distributed policy learning. This enables UAVs to collaboratively reinforce successful defense behaviors and adaptively coordinate MTD actions in real time. Compared to other multi-agent methods such as DESOLATER and ID-HAM, this approach ensures that high-performing agents have greater influence on the global policy, leading to more robust and consistent defense against dynamic and adaptive attacks.
@@
V-D Energy Consumption and Defense Cost Analysis
To estimate the energy consumption of the UAV swarm when equipped with different defense methods, we adopt a simplified model introduced by Liu et al. [44] to express the power of a UAV according to its velocity, which includes three components, namely induced power, profile power, and parasite power as
$P(v_{i})=(c_{1}+c_{2})\cdot(mg)^{3/2}+c_{3}\cdot v_{i}^{3}$ (31)
We can also observe that the energy consumption of DESOLATER is very close to our method under node attacks, which can be attributed to the fact that DESOLATER also employs a policy gradient-based approach. However, in the link attack scenario, PG-FMADRL enables more effective agent coordination by leveraging the federated learning framework. Consequently, under greedy attack strategies, PG-FMADRL achieves a 10.9% reduction in energy consumption compared to DESOLATER, further highlighting its effectiveness in collaborative defense and energy efficiency.
@@ V-D Energy Consumption and Defense Cost Analysis
In addition to energy consumption, we also analyze the cumulative defense cost of the UAV swarm, which is defined as the total cost associated with MTD execution during the entire episode. As we can see in Fig. 6, the cumulative defense cost varies significantly across different methods and attack scenarios. It is evident that the proposed method consistently achieves the lowest cumulative cost, regardless of the attack type or attacker strategy. This demonstrates the superior efficiency of our approach in minimizing unnecessary defense actions while maintaining robust protection. Furthermore, baseline methods such as WF-MTD and RE-MTD incur significantly increasing costs, as they tend to trigger more frequent or redundant actions due to their lack of adaptive coordination. In contrast, for both node and link attacks from fixed and random attackers, PG-FMADRL maintains a nearly flat cost curve, indicating that the swarm can effectively mitigate attacks with minimal resource expenditure. Specifically, PG-FMADRL achieves a cumulative cost of 1.16 under fixed link attacks when the episode ends, having a significant reduction of 98.3% compared to RE-MTD.
@@
The lower energy consumption and defense cost achieved by PG-FMADRL are attributed to its intelligent and context-aware action selection. By leveraging policy gradient optimization, each UAV can precisely determine when and which MTD mechanism to trigger, thus minimizing unnecessary or redundant actions. This fine-grained control ensures efficient use of resources while maintaining strong protection, outperforming baseline methods that often rely on more frequent or less targeted defense operations.
V-E Scalability Analysis

### v4

@@ V-A Simulation Setup
Federated aggregation interval $K_{\text{agg}}$ 20
@@ V-C Effectiveness and Efficiency Analysis
To analyze the defense effectiveness and operational efficiency of our method, we first evaluate the attack mitigation rate under various scenarios. We then assess the system’s resilience by measuring the recovery time required to restore connectivity after an attack. We compare our PG-FMADRL method against four representative MTD-based approaches, including WF-MTD, RE-MTD, ID-HAM, and DESOLATER.
@@
We can see that PG-FMADRL consistently achieves the highest or near-highest attack mitigation rates across all attack scenarios. For node attack scenario, PG-FMADRL achieves highest attack mitigation rates of 0.9975 and 0.9999 against fixed and random attackers, respectively, while maintaining the lowest standard deviation among all compared methods. This indicates superior stability and robustness. In the link attack scenario, although the attack mitigation rates show a slight decline, the proposed method still achieves the highest values, reaching 0.9969 for fixed and 0.9782 for random attackers. These results demonstrate the method’s strong and consistent defense effectiveness across different scenarios.
Notably, PG-FMADRL’s advantage becomes most apparent against greedy attackers. While ID-HAM and DESOLATER rely on independent learning where each agent must discover effective strategies in isolation, our federated architecture acts as a knowledge accelerator. For instance, in the challenging greedy link attack scenario, PG-FMADRL reaches an attack mitigation rate of 0.9367, significantly outperforming the other methods, which achieve only 0.7910, 0.7625, 0.7360, and 0.7721, respectively. This is because our reward-weighted aggregation mechanism allows the entire swarm to rapidly learn from the experiences of the most successful agents those achieving high rewards, by effectively countering the attack. Our collaborative learning ensures that effective defense policies are quickly propagated and adopted across the swarm, leading to a more robust and coordinated response.
Beyond mitigation rate, we also assess the efficiency of each method using recovery time, defined as the duration required for the system to restore normal connectivity following a disruption. Fig. 4 presents the recovery times of different defense methods.
As we can see, in the node attack scenario, WF-MTD shows the longest recovery times under fixed and random attackers, reaching 6.2 and 7.6 seconds, respectively. Under greedy attacks, RE-MTD performs the worst, with a recovery time of 18.7 seconds. In contrast, our PG-FMADRL method dramatically reduces the recovery time by approximately 94.6%, requiring about 1 second to restore a secure state. In the link attack scenario, the worst recovery times are observed for different methods depending on the attacker type. Such prolonged recovery durations leave the UAV swarm vulnerable for extended periods, increasing both operational risk and energy consumption. Although DESOLATER performs relatively well under greedy attacks in the node scenario, its recovery time increases significantly to 11.8 seconds in the link scenario. On the other hand, PG-FMADRL achieves the shortest recovery times across all attack scenarios and attacker types, enabling rapid mitigation and minimizing disruption.
In summary, the superior attack mitigation rates and reduced recovery times of PG-FMADRL stem from its reward-weighted aggregation and distributed policy learning. This enables UAVs to collaboratively reinforce successful defense behaviors in real time. Compared to other multi-agent methods such as DESOLATER and ID-HAM, this approach ensures that high-performing agents have greater influence on the global policy, leading to more robust and consistent defense against dynamic and adaptive attacks.
@@
V-D Energy Consumption and Defense Cost Analysis
@@
While the previous section established the effectiveness and efficiency of our method, a comprehensive evaluation must also consider the associated overhead. This section analyzes the operational costs from two perspectives, including the energy consumed by the UAVs during defensive maneuvers and the cumulative cost of executing MTD actions.
First, we adopt a simplified model introduced by Liu et al. [43] to express the power of a UAV according to its velocity, which includes three components, namely induced power, profile power, and parasite power as
$P(v_{i})=(c_{1}+c_{2})\cdot(mg)^{3/2}+c_{3}\cdot v_{i}^{3}$ (31)
We can also observe that the energy consumption of DESOLATER is very close to our method under node attacks, which can be attributed to the fact that DESOLATER also employs a policy gradient-based approach. However, our methods enables more effective agent coordination by leveraging the federated learning framework. Consequently, under greedy attack strategies, PG-FMADRL achieves a 10.9% reduction in energy consumption compared to DESOLATER, highlighting its effectiveness in collaborative defense and energy efficiency.
In addition to energy consumption, we analyze the cumulative defense cost, which reflects the resource expenditure from triggering MTD mechanisms. As we can see in Fig. 6, the cumulative defense cost varies significantly across different methods and attack scenarios. It is evident that the proposed method consistently achieves the lowest cumulative cost, regardless of the attack type or attacker strategy. This demonstrates the superior efficiency of our approach in minimizing unnecessary defense actions while maintaining robust protection. Furthermore, baseline methods such as WF-MTD and RE-MTD incur significantly increasing costs, as they tend to trigger more frequent or redundant actions due to their lack of adaptive coordination. In contrast, for both node and link attacks from fixed and random attackers, PG-FMADRL maintains a nearly flat cost curve, indicating that the swarm can effectively mitigate attacks with minimal resource expenditure. Specifically, PG-FMADRL achieves a cumulative cost of 1.16 under fixed link attacks when the episode ends, having a significant reduction of 98.3% compared to RE-MTD.
@@
The lower energy consumption and defense cost achieved by PG-FMADRL are attributed to its intelligent and context-aware action selection. By leveraging policy gradient optimization, each UAV can precisely determine when and which MTD mechanism to trigger, thus minimizing unnecessary actions. This fine-grained control ensures resource efficiency while maintaining strong protection, outperforming baseline methods that rely on more frequent or less targeted defense operations.
V-E Scalability Analysis